OpenStudy (kayy):

Multipy the following polynomials... (8x + 3)(4x + 7)

OpenStudy (anonymous):

use Distributive Property,. Do u know it???

OpenStudy (kayy):

P-E-M-D-A-S???

OpenStudy (anonymous):

no, I think in your country it is called FILA or something.... (a+b) (c+d) = a(c+d) +b(c+d) = ac + ad +bc +bd

OpenStudy (amistre64):

no need to do distributive; just multiply like normal ... 8x + 3 4x + 7 ------------- 56x +21 32x^2 +12x ---------------- ?? ?? ??

OpenStudy (anonymous):

Yeah it is called FOIL....

OpenStudy (kayy):

Ok, yeah. I've used FOIL before.